There are 1/5 as many rides as the games at the fair if the total number of is 20 how many rides are there
Answer by Alan3354(69443)

There are 1/5 as many rides as the games at the fair if the total number of is 20 how many rides are there
---------------------
"total number of ???? "
------
G = 5R
G + R = 20
-----
5R + R = 20
R = 10/3 = 3 1/3 rides.
-------------------------
1/3 of a ride should not be expensive.